OpenStudy (anonymous):

Scientists in a laboratory produce a sound wave in a simulated environment (air) with a velocity of 419 m/s. What is the temperature (in Kelvin) of the environment?

OpenStudy (rajat97):

We can assume speed of sound in air to be \[\sqrt{constant \times T}\] where T is the temperature of air. we know, at 20 deg C(293K), the speed of sound in air is 343.2m/s so we can use unitary method to solve this. \[\sqrt{293}\] corresponds to 343.2m/s so \[\sqrt{T}\] correspond to 419m/s so we get \[\sqrt{T}\]=20.898 thus, T=436.726K i.e. 153.726 deg C

OpenStudy (rajat97):

i hope this helps you